# Data Release 1 of the Dark Energy Spectroscopic Instrument

### v1

@@ F.5 Lyman-Alpha Forest VACs
Lyman Alpha Forest Year 1 Deltas VAC contains the flux-transmission field used in the DR1 Ly BAO measurement. The fluctuations in the calibration region (rest-frame wavelength from 1600 to 1850 Å), the Ly region A (rest-frame wavelength from 1040 to 1205 Å), and Ly region B (rest-frame wavelength from 920 to 1020 Å) were obtained using Python Code for IGM Cosmological-Correlations Analyses (PICCA; du Mas des Bourboux et al. 2020). The continuum-fitting process is discussed in detail by Ramírez-Pérez et al. (2024). This catalog provides the estimated quasar continua, the fluctuations in the spectra, and their associated weights for all lines-of-sight for each HEALPix pixel. The fitted variance functions, stack of fluctuations, and the mean continuum are stored for each iteration during the continuum-fitting process.
Lyman Alpha Forest Year 1 Correlations VAC contains the correlation data products that are used in the DR1 Ly BAO measurement. It contains correlation measurements between Ly forest fluctuations in both regions A (1040 to 1205 Å) and B (920 to 1020 Å), and tracer quasars. For each of the four combinations, we provide the correlation function and the distortion matrix. The latter is used to forward-model the effect of reconstructing the unabsorbed flux of quasar spectra (see, e.g., du Mas des Bourboux et al. 2020). Furthermore, we provide one smoothed covariance matrix that includes the cross-covariances between the different correlation function measurements.

### v2

@@ F.5 Lyman-Alpha Forest VACs
Lyman Alpha Forest Year 1 Deltas VAC contains the flux-transmission field used in the DR1 Ly BAO measurement. The fluctuations in the calibration region (rest-frame wavelength from 1600 to 1850 Å), the Ly region A (rest-frame wavelength from 1040 to 1205 Å), and Ly region B (rest-frame wavelength from 920 to 1020 Å) were obtained using Python Code for IGM Cosmological-Correlations Analyses (PICCA; du-mas-des-bourboux20a). The continuum-fitting process is discussed in detail by ramirez-perez24a. This catalog provides the estimated quasar continua, the fluctuations in the spectra, and their associated weights for all lines-of-sight for each HEALPix pixel. The fitted variance functions, stack of fluctuations, and the mean continuum are stored for each iteration during the continuum-fitting process.
Lyman Alpha Forest Year 1 Correlations VAC contains the correlation data products that are used in the DR1 Ly BAO measurement. It contains correlation measurements between Ly forest fluctuations in both regions A (1040 to 1205 Å) and B (920 to 1020 Å), and tracer quasars. For each of the four combinations, we provide the correlation function and the distortion matrix. The latter is used to forward-model the effect of reconstructing the unabsorbed flux of quasar spectra (see, e.g., du-mas-des-bourboux20a). Furthermore, we provide one smoothed covariance matrix that includes the cross-covariances between the different correlation function measurements.